OpenStudy (saifoo.khan):

explain the characteristics and parts of a logarithmic function and its graph.

OpenStudy (anonymous):

sorry that's inversely exponentially over my head

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

that's Alg2 man.

OpenStudy (jamesj):

Well, let's see. The domain of ln is (0,infinity) and has a range of the entire real line; it's strictly increasing; has one root at x = 1; has a vertical asymptote as x->0+ of the negative y axis. What else?

OpenStudy (jamesj):

It has negative second derivative everywhere which means that while it is increasing, it increases at a slowing rate.

OpenStudy (saifoo.khan):

Q2: explain the similarities and differences between the graphs of a radical function and a logarithmic function. 

OpenStudy (jamesj):

f(x) = 1/x would be a radical function, for instance?

OpenStudy (saifoo.khan):

no it wont be.

OpenStudy (jamesj):

what's the definition you are using for a radical function?

OpenStudy (saifoo.khan):

the one with radicals. LOL.

OpenStudy (saifoo.khan):

\[f(x) = \sqrt{x} -2\]Something like that.

OpenStudy (jamesj):

I see. Well, often radical function also have restricted domains like the ln function; i.e., not the whole real line. O

OpenStudy (jamesj):

Often radical functions also have second derivative of just one sign and first derivative also of one sign. But without a sharper definition of a radical function I'm finding it hard to characterize this in general.

OpenStudy (saifoo.khan):

Log functions also have restrictions.

OpenStudy (jamesj):

for example f(x) = sqrt(x) is only defined on [0,infty), almost like the ln function; it is tricly increasing and has negative second derivative everywhere. Has only one root on its domain. But unlike the ln function its range is [0,infty)

OpenStudy (jamesj):

that'll probably do.

OpenStudy (saifoo.khan):

Alright. thanks.